OpenStudy (anonymous):

I was doing a probability and statistics worksheet and I got stuck on a basic mathematical thing involving factorials. Could anyone explain why x!(x+1) is the same as (x+1)! ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

lets pick n = 6 then \(6!=6\times 5\times 4\times 3\times 2\times 1\) and \(n+1=7\) now \(7!=7\times 6\times 5\times 4\times 3\times 2\times 1=7\times 6!\)

OpenStudy (anonymous):

yes I worked that out using wolfram. But the thing is, in an exam how would I realise this. Is there a specific method to simplifying factorials or is it just by intuition?

OpenStudy (anonymous):

i hate to say that it is obvious, but it is obvious consider any number n then \((n+1)!=(n+1)\times n\times (n-1)\times ...\times 2\times 1\) and this is fairly clearly the same as \[(n+1)\times n!\]

OpenStudy (anonymous):

sometimes factorials are even defined that way \[1!=1, n!=n(n-1)!\]

OpenStudy (anonymous):

you might also notice that \(\frac{(n+1)!}{n!}=n+1\) for example

OpenStudy (anonymous):

I know these look obvious but I cant see it. lol

OpenStudy (anonymous):

really try it with numbers and it will be clear

OpenStudy (anonymous):

like the n(n-1)! = n! I dont see them as equal unless I start plugging in values using a calculator

OpenStudy (anonymous):

how to you get 4! ? by multiplying 4*3*2*1 how do you get 5! ? by multiplying 5*4*3*2*1

OpenStudy (anonymous):

so they are the same except for the 5

OpenStudy (anonymous):

so basically there is no specific method of simplifying ?

OpenStudy (anonymous):

don't think about what the final answer is, just think about what you are computing

OpenStudy (anonymous):

\[\frac{7!}{5!}=\frac{7\times 6\times 5\times 4\times 3\times 2\times 1}{5\times 4\times 3\times 2\times 1}=7\times 6\]

OpenStudy (anonymous):

with a little practice you would know right away that \(\frac{7!}{5!}=7\times 6\)

OpenStudy (anonymous):

ok. Think i'm more confident with these now with your help. Just need more practice

OpenStudy (anonymous):

Thanks for your help :)

OpenStudy (anonymous):

yw